OpenStudy (brekathleen):

Simplify the following expression : 8x-10y+5z-5x-16+6y a) 3x-6y+11z b) 13x-6y+11z c) 13x-26y+11z D)3x-26y+11z

OpenStudy (er.mohd.amir):

add like terms to each others

OpenStudy (brekathleen):

how though ?

OpenStudy (er.mohd.amir):

8x-5x=?

OpenStudy (mathstudent55):

Like terms are terms that have the same variables. For example, in 5x + 2y + 6x - 7y, there are two sets of like terms. The terms 5x and 6x are like terms. The terms 2y and -7y are also like terms. You can move the like terms together, then you combine them. \(5x + 2y + 6x - 7y\) \(= 5x + 6x + 2y - 7y\) \(= 11x - 5y\) That is the final answer. Since 11x and -5y are not like terms, they can't be combined.

OpenStudy (mathmale):

Examples: 8x and 5x are "like terms" and can be combined. 8x-5x=3x. 8x and 5y are not like terms and must be left as they are.